# Collecting Cost-Effective, High-Quality Truthfulness Assessments with LLM Summarized Evidence

### v1

@@ Abstract
With the degradation of guardrails against mis- and disinformation online, it is more critical than ever to be able to effectively combat it. In this paper, we explore the efficiency and effectiveness of using crowd-sourced truthfulness assessments based on condensed, large language model (LLM) generated summaries of online sources. We compare the use of generated summaries to the use of original web pages in an A/B testing setting, where we employ a large and diverse pool of crowd-workers to perform the truthfulness assessment. We evaluate the quality of assessments, the efficiency with which assessments are performed, and the behavior and engagement of participants. Our results demonstrate that the Summary modality, which relies on summarized evidence, offers no significant change in assessment accuracy over the Standard modality, while significantly increasing the speed with which assessments are performed. Workers using summarized evidence produce a significantly higher number of assessments in the same time frame, reducing the cost needed to acquire truthfulness assessments. Additionally, the Summary modality maximizes both the inter-annotator agreements as well as the reliance on and perceived usefulness of evidence, demonstrating the utility of summarized evidence without sacrificing the quality of assessments.
@@
The proliferation of online misinformation has made truthfulness assessment a crucial task in the digital era. Fact-checking efforts, traditionally conducted by experts, have faced challenges in scalability and timeliness as the volume of information continues to grow [5]. Recent developments, such as X and Meta scaling down their reliance on paid fact-checkers, [^1] https://www.nytimes.com/2025/01/07/business/meta-community-notes-x.html. further highlight the urgency of scalable and efficient truthfulness assessment methods. As social media platforms reduce investments in professional fact-checking, alternative approaches become increasingly vital for addressing the growing volume of online misinformation. To address this, crowdsourcing has emerged as a promising alternative, leveraging the collective efforts of non-expert workers to evaluate truthfulness efficiently [44, 33]. However, the success of crowdsourced truthfulness assessments relies heavily on the format, presentation, and quality of evidence provided to participants. The choice of evidence itself, along with how it is presented, directly influences workers’ ability to make accurate and reliable judgments.
@@
In this work, we focus on comparing two approaches: a Standard modality, where participants evaluate full-length webpages, and a Summary modality, where participants assess summarized evidence generated by a state-of-the-art LLM. The summarized format aims to reduce cognitive load and improve worker efficiency by condensing essential information while preserving factual accuracy. This presentation allows participants to process more information in less time, increasing overall throughput. Such improvements not only offer productivity gains but also translate into significant cost savings for large-scale annotation tasks, making summarization particularly valuable. However, the effectiveness and trade-offs of these two modalities remain an open question, which we address through a detailed empirical comparison using an A/B testing framework. This experimental design allows us to systematically compare the two modalities under controlled conditions, providing insights into how evidence presentation affects truthfulness evaluations. To guide our analysis, we address the following Research Questions:
@@ 3. RQ3:
Multiple studies have explored the application of crowdsourcing for misinformation detection, demonstrating its potential as a scalable alternative to expert fact-checking [27, 38]. La Barbera et al. [26] highlighted the influence of worker bias in assessing statements, an issue further explored by Roitero et al. [43], who observed similar agreement levels across various truthfulness scales. Expanding on this, Soprano et al. [48] and Liu et al. [30] introduced multidimensional truthfulness scales, showing that crowd judgments are reliable across multiple truthfulness dimensions, each capturing different aspects of truthfulness. The comparative performance of crowd workers and experts has also been examined. For example, Zhao and Naaman [55] found that while the crowd performed on par with experts in terms of accuracy, objectivity, and clarity, their assessments often depended on existing professional knowledge. Similarly, Saeed et al. [45] analyzed social network data and demonstrated that crowd evaluations relied on different evidence sources compared to experts, with better scalability and efficiency. Allen et al. [3] further highlighted the scalability of crowd-based methods, noting that political alignment influenced crowd judgments. Additionally, psychological [39] and linguistic [51] peculiarities in misinformation were explored, emphasizing the diverse factors affecting the quality of crowdsourced truthfulness evaluations.
@@
Although this study focuses on crowdsourcing, advances in automatic and hybrid fact-checking methods have contributed significantly to the field. Fully automated techniques, including machine learning and deep learning models, have been widely explored for misinformation detection [2, 5, 9, 22, 32]. These methods have been evaluated using public datasets and benchmarks provided by initiatives like CLEF CheckThat! Lab [34, 35] and studies such as Augenstein et al. [6]. Comparative analyses also have been developed to study variability in model performance across datasets [22, 41, 46, 49]. Recent research has also focused on generating human-readable explanations [4, 7, 50, 53] and incorporating complex interactions within machine learning pipelines [29]. Hybrid approaches that integrate human input with machine learning have gained attention as well [10, 40, 52].
In this paper, we leverage automatic signals generated by a state-of-the-art LLM to enhance the efficiency and scalability of a crowd-based approach for truthfulness assessment. By combining LLM-generated summaries with crowdsourced evaluations, we aim to investigate the interplay between automated summarization and human judgment in achieving reliable and cost-effective misinformation detection.
@@
In this section, we detail the methodology employed in our experiment. To promote transparency and reproducibility, all data collected during the experiment are made publicly available to the research community.222(anonymized link for peer-review) https://osf.io/bnfxm/?view_only=9dcfe098a1ea440b89adbc4574838ff9.
@@
This study utilizes the same dataset and adopts a similar experimental design as reported by La Barbera et al. [25], with modifications tailored to meet the specific objectives and requirements of our research. To ensure this paper is self-contained and understandable independently, we report both the details of the data and crowdsourcing task employed. The dataset we use is composed of a subset of statements sourced from the PolitiFact website, [^3] https://www.politifact.com/ as also done by previous research [11, 25, 48]. PolitiFact has been an active platform for fact-checking statements from U.S. politicians, political organizations, public figures, and posts on social media since 2007, accumulating over 24,000 fact-checks with regular updates.
PolitiFact statements are categorized by expert judges using a six-level truthfulness scale having levels: pants-on-fire, false, mostly-false, half-true, mostly-true, and true. In our study, we adopt this same six-level scale to maintain consistency and enable direct comparisons between the expert fact-checkers and the crowdsourced assessments. As done in previous work, we select a subset of 120 statements (20 for each ground truth category) from PolitiFact. Additionally, we made sure each was supported by at least 10 webpages retrieved as evidence by using the Bing Web Search API. [^4] https://www.microsoft.com/en-us/bing/apis/bing-web-search-api These statements form the basis of the crowdsourcing experiments detailed in the subsequent sections.
@@
Table 1: Example of LLM generated summary given the query ‘Opposition to having a fully elected Chicago Board of Education is in the “super minority.”‘.
@@
We process the evidence related to each statement to create two distinct versions for evaluation: the original, unaltered full-length web page and a more concise summarized version, obtained by leveraging a LLM. To achieve the summarization version, we employ the capabilities of the Meta-Llama-3-8BInstruct model, a state-of-the-art model known for its robust performance in natural language understanding and generation tasks [13]. [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
The process begins with a carefully crafted prompt that we provide to the model. This prompt is specifically designed to guide the model to extract and summarize the essential factual elements from the content of a webpage, ensuring that the generated summary remains the same with respect to the original information in terms of factuality and stance. After several rounds of trial and error, the exact wording of the prompt we use is given in Figure 1.
@@
We adopt the crowdsourcing task as detailed by La Barbera et al. [25], modifying it to collect questions for our study while preserving the core structure of the original design. This design, originally developed and validated in earlier research [11, 25, 48], allows us to collect truthfulness assessments at scale using crowdsourcing.
@@
After filling a demographic questionnaire, the participants perform our task using a structured interface where they first encounter the statement to be assessed, followed by a set of evidences. Each worker sees the title of the web page along with a snippet, similar to the presentation layout of a search engine. Then, depending on the task layout, if the worker clicks on the link they see either the full webpage or its summary. Subsequent to evidence review and selection, participants assess the statement. We ask workers to first evaluate the overall truthfulness of the statement using the same six-level scale as the one used by experts. Then, we ask the workers to answer two questions regarding how useful they found the evidence in determining the truthfulness of the statement: “How useful was the evidence?” and “Did you use the evidence to fact-check the statement, or did you have a clear opinion before reviewing the evidence?”. For both questions, we collect responses on a five-level Likert scale. The scale for the first question includes the following options: “Not Useful at All”, “Slightly Useful”, “Moderately Useful”, “Very Useful”, and “Extremely Useful”. The scale for the second question includes: “I did not consider the evidence at all before forming my opinion”, “I considered the evidence slightly before forming my opinion”, “I considered the evidence as much as my initial opinion before forming my final opinion”, “I considered the evidence more than my initial opinion before forming my final opinion”, and “I fully based my opinion on the evidence after reviewing it”. To ensure high-quality assessments from workers, we implement the following checks: a minimum time requirement of 3 seconds per statement and correct evaluations of two gold-standard questions, one clearly true and the other clearly false.
@@
The answers to the demographic questionnaire allow us to assess the demographics of participants across the two experimental conditions. Table 2 presents the distribution of political considerations and views for each condition. The results indicate a balanced representation across both modalities, with comparable percentages for both the political affiliation and ideological orientation. We also measure abandonment [21] to evaluate participant engagement. Initial abandonment (49% for standard, 38% for summary) and mid-task abandonment (10% for standard, 13% for summary) are comparable, with no statistically significant differences observed.
@@
Figure 2: Agreement on individual and aggregated assessments between workers and experts (here and in the following figures we abbreviate the six Politifact levels with their initials).
@@
We inspect the agreement between workers and experts, shown in Figure 2. We discuss both the individual and the aggregated scores for both the Standard and Summary modalities. In terms of individual scores, in the Standard modality, we observe an accuracy of 0.31, a MSE of 3.06, and a MAE score of 1.27. The Summary modality reports similar accuracy (0.31) but exhibits a slightly higher MSE (3.27) and MAE (1.30). This similarity in the metric scores suggest comparable quality level between the standard and summary modalities, suggesting that the summarized content does not impact workers’ effectiveness, which is not trivial as the content presented to the participants in the summary version of the task is substantially condensed. We further examined whether the observed differences in MAE and MSE between the modalities are statistically significant. Using bootstrapped confidence intervals [^6] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.bootstrap.html. with 10,000 resamples, we estimated the range of plausible values for the difference in these metrics. For MAE, the mean difference was $-0.028$ with a 95% confidence interval ranging from $-0.167$ to $0.112$. Similarly, for MSE, the mean difference was $-0.212$ with a 95% confidence interval spanning from $-0.767$ to $0.340$. As both intervals include zero, we conclude that there is no statistically significant difference between the Standard and Summary modalities in terms of MAE and MSE.
@@
We now turn to aggregated scores, still shown in Figure 2. The scores are aggregated using three distinct statistical methods: mean, mode (i.e., majority voting), and median. The mean offers a general view of central tendency and has proven to lead to higher agreement with expert assessments [25, 48], while the mode identifies the most frequently occurring outcome, and the median provides an additional robust measure of central tendency that is less sensitive to extreme values. In the Standard modality, mean accuracy is slightly higher ($0.33$) compared to the Summary modality ($0.31$), though the difference is minimal. Similarly, the Standard modality shows marginally better error metrics, with MSE ($1.65$ vs. $1.77$) and MAE ($0.95$ vs. $1.00$) indicating lower dispersion and deviation from the ground truth. Mode aggregation reflects a similar trend, with accuracy at $0.42$ for the Standard modality and $0.38$ for the Summary modality, alongside smaller errors (MSE: $1.81$ vs. $2.13$, MAE: $0.91$ vs. $0.98$). Median aggregation further supports this finding, with the Standard modality showing slightly better performance across metrics (accuracy: $0.40$ vs. $0.38$, MSE: $1.80$ vs. $2.13$, MAE: $0.92$ vs. $0.93$). While errors in the Summary modality are slightly more pronounced, both approaches achieve comparable accuracy levels. Bootstrap analysis confirms no statistically significant differences across accuracy and error measures. Overall, both modalities show similar effectiveness across several metrics, indicating that, despite the reduced volume of information provided to workers in the Summary modality, it closely resembles the Standard modality in terms of data quality. To test whether worker demographics influence the effectiveness of truthfulness assessments, we additionally run an Ordinary Least Squares (OLS) regression analysis [19]. The results show no statistically significant relationship between demographics and accuracy, indicating that these factors do not substantially impact assessment performance.
@@
Figure 3 shows the values of internal agreement across the two task modalities, both for the whole task (dotted lines) as well as for the single ground truth levels of the statements. For the Standard modality, the $\alpha$ values are generally lower than the Summary modality, reflecting overall less agreement among participants, with the only exception of the pants-on-fire truthfulness level. This suggests that when participants are presented with full-length content, there may be a greater diversity in how the information is interpreted, possibly due to the complexity and volume of the information provided. Furthermore, the lowest $\alpha$ values appear in the categories in the middle of the scale, indicating that the claims with no extreme truthfulness score are also the most challenging for participants to consistently evaluate. Conversely, the Summary modality shows higher $\alpha$ values across all categories except pants-on-fire and for the overall task, which is also statistically significant at the $p<0.05$ level according to the Mann-Whitney U Test, indicating a higher degree of consistency among workers ratings. This improvement in internal agreement could be attributed to the condensed nature of the content, which might focus the workers evaluations and reduce both the cognitive load as well as the ambiguity or variability in interpretation that more extensive evidence might introduce. Notably, the alpha values are particularly higher in categories like mostly-false and true where the quality of the summarization may help clarify the key aspects of the statements.
Overall, these findings remark the impact of evidence summarization on the consistency of workers assessments. The higher internal agreement in the Summary modality suggests that summarization not only maintains the essential content needed for the truthfulness assessment but may also aid in aligning worker scores, leading to more uniform interpretations and assessments of the veracity of statements.
@@
Table 3: Error types across truthfulness categories.
@@
Standard pants-on-fire 24 (24.0%) 76 (76.0%)
@@
mostly-false 15 (18.3%) 51 (62.2%) 34 (19.5%)
half-true 33 (48.5%) 41 (30.1%) 26 (20.6%)
mostly-true 51 (50.0%) 15 (14.7%) 34 (33.3%)
@@
Summary pants-on-fire 26 (26.0%) 74 (74.0%)
@@
mostly-false 12 (15.6%) 47 (61.0%) 41 (23.4%)
half-true 36 (42.9%) 34 (40.5%) 13 (16.6%)
mostly-true 40 (50.6%) 28 (32.0%) 32 (17.4%)
@@
A more fine-grained analysis, shown in the second part of Table 3, highlights the distribution of errors across modalities. In the Standard modality, overestimation emerges as the most frequent error type for categories such as mostly-false (62.2%). Conversely, correct assessments are most common in categories like mostly-true (50.0%) and half-true (48.5%), suggesting workers find it easier to correctly evaluate statements closer to the middle of the truthfulness spectrum. Underestimation occurs less frequently overall but is more pronounced for higher truthfulness categories. In the Summary modality, overestimation remains dominant in the same categories as for the Standard modality. However, correct assessments are slightly more evenly distributed, with the highest rates observed for mostly-true (50.6%) and half-true (42.9%). Overall, these findings suggest that both modalities exhibit similar patterns with respect to error types. We further analyze the mismatch magnitude for overestimation and underestimation errors across the Standard and Summary modalities and find similar error patterns. Both modalities exhibit comparable average mismatch values for overestimation (1.99) and slightly more negative values for underestimation in the Summary modality (-1.73 vs. -1.65). None of these differences are statistically significant, confirming the comparability of the two approaches.
@@
The results, shown in Figure 4, reveal a trend where the accuracy of both modalities slightly improves (although no statistical significance is observed) with a lower variance as the number of workers per document increases. This pattern is consistent with expectations from crowdsourcing settings, where aggregating multiple assessments tends to mitigate individual biases and errors. Notably, the Summary modality achieves effectiveness levels comparable to the Standard modality across all sampling sizes. These indicate that summarized evidence can provide comparable results to full-length evidence in terms of both accuracy and consistency, and that this equivalence is achieved while potentially reducing the cognitive load on workers and the resources required for task completion.

### v2

# Efficiency and Effectiveness of LLM-Based Summarization of Evidence in Crowdsourced Fact-Checking

@@ Abstract
Evaluating the truthfulness of online content is critical for combating misinformation. This study examines the efficiency and effectiveness of crowdsourced truthfulness assessments through a comparative analysis of two approaches: one involving full-length webpages as evidence for each claim, and another using summaries for each evidence document generated with a LLM. Using an A/B testing setting, we engage a diverse pool of participants tasked with evaluating the truthfulness of statements under these conditions.
Our analysis explores both the quality of assessments and the behavioral patterns of participants. The results reveal that relying on summarized evidence offers comparable accuracy and error metrics to the standard modality while significantly improving efficiency. Workers in the Summary setting complete a significantly higher number of assessments, reducing task duration and costs. Additionally, the Summary modality maximizes internal agreement and maintains consistent reliance on and perceived usefulness of evidence, demonstrating its potential to streamline large-scale truthfulness evaluations.
Keywords: Truthfulness Assessment, Summarization, Large Language Models
@@
The proliferation of online misinformation has made truthfulness assessment a crucial task in the digital era. Fact-checking efforts, traditionally conducted by experts, have faced challenges in scalability and timeliness as the volume of information continues to grow [6]. Recent developments, such as X and Meta scaling down their reliance on paid fact-checkers, [^1] https://www.nytimes.com/2025/01/07/business/meta-community-notes-x.html. further highlight the urgency of scalable and efficient truthfulness assessment methods [8]. As social media platforms reduce investments in professional fact-checking, alternative approaches become increasingly vital for addressing the growing volume of misinformation. To address this, crowdsourcing has emerged as a promising alternative, leveraging the collective efforts of non-expert workers to evaluate truthfulness efficiently [46, 36]. However, the success of crowdsourced truthfulness assessments relies heavily on the format, presentation, and quality of evidence provided to participants. The choice of evidence itself, along with how it is presented, directly influences workers’ ability to make accurate and reliable judgments.
@@
In this work, we focus on comparing two approaches: a Standard modality, in which participants evaluate full-length webpages, and a Summary modality, in which participants assess summarized evidence generated by a state-of-the-art LLM, Meta-Llama-3-8BInstruct [16]. The summarized format aims to reduce cognitive load and improve worker efficiency by condensing essential information while preserving factual accuracy. This presentation allows participants to process more information in less time, increasing overall throughput. Such improvements not only offer productivity gains but also translate into significant cost savings for large-scale annotation tasks, making summarization particularly valuable. However, the effectiveness and trade-offs of these two modalities remain an open question, which we address through a detailed empirical comparison using an A/B testing framework. This experimental design allows us to systematically compare the two modalities under controlled conditions, providing insights into how evidence presentation affects truthfulness evaluations. To guide our analysis, we address the following Research Questions (RQs):
@@ 3. RQ3:
Multiple studies have explored the application of crowdsourcing for misinformation detection, demonstrating its potential as a scalable alternative to expert fact-checking [30, 41]. La Barbera et al. [29] highlighted the influence of worker bias in assessing statements, an issue further explored by Roitero et al. [46], who observed similar agreement levels across various truthfulness scales. Expanding on this, Soprano et al. [51] and Liu et al. [33] introduced multidimensional truthfulness scales, showing that crowd judgments are reliable across multiple truthfulness dimensions, each capturing different aspects of truthfulness. The comparative performance of crowd workers and experts has also been examined. For example, Zhao and Naaman [62] found that while the crowd performed on par with experts in terms of accuracy, objectivity, and clarity, their assessments often depended on existing professional knowledge. Similarly, Saeed et al. [47] analyzed social network data and demonstrated that crowd evaluations relied on different evidence sources compared to experts, with better scalability and efficiency. Allen et al. [3] further highlighted the scalability of crowd-based methods, noting that political alignment influenced crowd judgments. Additionally, linguistic [56] and psychological [42] peculiarities in misinformation were explored, emphasizing the diverse factors affecting the quality of crowdsourced truthfulness evaluations.
@@
Although this study focuses on crowdsourcing, advances in automatic and hybrid fact-checking methods have contributed significantly to the field. Fully automated techniques, mostly based on machine learning models, have been widely explored for misinformation detection [6, 12, 2, 11, 35, 25, 5]. These methods have been evaluated using public datasets and benchmarks, including those provided by initiatives like the CLEF CheckThat! Lab [37, 38] as well as other projects such as MultiFC [7], FEVER [53, 54], AVeriTeC [48] and Factcheck-Bench [57]. Comparative analyses also have been developed to study variability in model performance across datasets [52, 25, 44, 49]. Recent research has also focused on generating human-readable explanations [4, 9, 60, 55] and incorporating complex interactions within machine learning pipelines [32]. Hybrid approaches that integrate human input with machine learning have gained attention as well [13, 59, 43]. Lastly, there is research on how fact checkers use tools in practice [58].
In this paper, we leverage automatic signals generated by a state-of-the-art LLM, Meta-Llama-3-8BInstruct [16], to enhance the efficiency and scalability of a crowd-based approach for truthfulness assessment. By combining LLM-generated summaries with crowdsourced evaluations, we aim to investigate the interplay between automated summarization and human judgment in achieving reliable and cost-effective misinformation detection.
@@
In this section, we detail the methodology employed in our experiment. To promote transparency and reproducibility, all data collected during the experiment are made publicly available to the research community. [^2] https://doi.org/10.17605/OSF.IO/BNFXM.
@@
This study uses the same dataset and adopts a similar experimental design to that reported by La Barbera et al. [28], with modifications tailored to meet the specific objectives and requirements of our research. To ensure that this paper is self-contained and understandable independently, we report both the details of the dataset and the crowdsourcing task employed. The dataset consists of a subset of statements sourced from the PolitiFact website, [^3] https://www.politifact.com/ as in previous research [51, 14, 28]. PolitiFact has been an active platform for fact-checking statements made by U.S. politicians, political organizations, public figures, and social media posts since 2007, accumulating over 24,000 fact-checks with regular updates.
PolitiFact statements are categorized by expert judges using a six-level truthfulness scale, consisting of the following levels: Pants On Fire, False, Mostly False, Half True, Mostly True, and True. In this study, we adopt the same six-level scale to maintain consistency and enable direct comparisons between expert fact-checkers and crowdsourced assessments. As in previous work, we select a subset of 120 statements (20 for each ground truth category) from PolitiFact. Additionally, we ensure that each statement is supported by at least 10 webpages retrieved as evidence using the Bing Web Search API. [^4] https://www.microsoft.com/en-us/bing/apis/bing-web-search-api These statements form the basis of the crowdsourcing experiments described in the subsequent sections.
@@
Table 1: Example of an LLM-generated summary given the query ‘Opposition to having a fully elected Chicago Board of Education is in the “super minority.”’.
@@
We process the evidence related to each statement to create two distinct versions for evaluation: the original, unaltered full-length web page and a more concise summarized version, obtained by leveraging a LLM. To achieve the summarization version, we employ the capabilities of the Meta-Llama-3-8B-Instruct model, a state-of-the-art model known for its robust performance in natural language understanding and generation tasks [16]. [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
The process begins with a carefully crafted prompt provided to the model. This prompt is specifically designed to guide the model in extracting and summarizing the essential factual elements from the content of a webpage, ensuring that the generated summary maintains consistency with the original information in terms of factuality and stance. After several rounds of trial and error, the exact wording of the prompt we use is given in Figure 1.
@@
We adopt the crowdsourcing task as detailed in La Barbera et al. [28], modifying it to collect responses for our study while preserving the core structure of the original design. This design, which was originally developed and validated in earlier research [28, 51, 14], allows us to collect truthfulness assessments at scale using crowdsourcing.
@@
After filling a demographic questionnaire, the participants perform our task using a structured interface where they first encounter the statement to be assessed, followed by a set of evidence. Each worker sees the title of the web page along with a snippet, similar to the presentation layout of a search engine. Then, depending on the task layout, if the worker clicks on the link they see either the full webpage or its summary. Subsequent to evidence review and selection, participants assess the statement. We ask workers to first evaluate the overall truthfulness of the statement using the same six-level scale as the one used by experts. Then, we ask the workers to answer two questions regarding how useful they found the evidence in determining the truthfulness of the statement: “How useful was the evidence?” and “Did you use the evidence to fact-check the statement, or did you have a clear opinion before reviewing the evidence?”. For both questions, we collect responses on a five-level Likert scale. The scale for the first question includes the following options: “Not Useful At All”, “Slightly Useful”, “Moderately Useful”, “Very Useful”, and “Extremely Useful”. The scale for the second question includes: “I did not consider the evidence at all before forming my opinion”, “I considered the evidence slightly before forming my opinion”, “I considered the evidence as much as my initial opinion before forming my final opinion”, “I considered the evidence more than my initial opinion before forming my final opinion”, and “I fully based my opinion on the evidence after reviewing it”. To ensure high-quality assessments from workers, we implement the following checks: a minimum time requirement of 3 seconds per statement and correct evaluations of two gold-standard questions, one clearly true and the other clearly false.
@@
The answers to the demographic questionnaire enable us to assess the demographics of participants across the two experimental conditions. Table 2 presents the distribution of political considerations and views for each condition. The results indicate a balanced representation across both modalities, with comparable percentages for both political affiliation and ideological orientation.
We also measure abandonment [24] to evaluate participant engagement. Initial abandonment rates (49% for Standard, 38% for Summary) and mid-task abandonment rates (10% for Standard, 13% for Summary) are comparable, with no statistically significant differences observed.
@@
Figure 2: Agreement between workers and experts on individual and aggregated assessments.
@@
We inspect the agreement between workers and experts, shown in Figure 2, where we report results for both individual and aggregated assessments. In all figures, the six PolitiFact levels are abbreviated with their initials. We discuss both the individual and the aggregated scores for the Standard and Summary modalities. In terms of individual scores, in the standard modality, we observe an accuracy of 0.31, a MSE of 3.06, and a MAE of 1.27. The summary modality reports a similar accuracy (0.31) but exhibits a slightly higher MSE (3.27) and MAE (1.30). This similarity in the metric scores suggests a comparable quality level between the standard and summary modalities, indicating that the summarized content does not impact workers’ effectiveness. This finding is nontrivial, given that the content presented in the summary version of the task is substantially condensed.
We further examined whether the observed differences in MAE and MSE between the modalities are statistically significant. Using bootstrapped confidence intervals [^6] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.bootstrap.html. with 10,000 resamples, we estimated the range of plausible values for the difference in these metrics. For MAE, the mean difference was $-0.028$ with a 95% confidence interval ranging from $-0.167$ to $0.112$. Similarly, for MSE, the mean difference was $-0.212$ with a 95% confidence interval spanning from $-0.767$ to $0.340$. As both intervals include zero, we conclude that there is no statistically significant difference between the standard and summary modalities in terms of MAE and MSE.
@@
We now turn to the aggregated scores, which are still shown in Figure 2. The scores are aggregated using three distinct statistical methods: mean, mode (i.e., majority voting), and median. The mean offers a general view of central tendency and has proven to lead to higher agreement with expert assessments [28, 51], while the mode identifies the most frequently occurring outcome, and the median provides an additional robust measure of central tendency that is less sensitive to extreme values. In the standard modality, mean accuracy is slightly higher ($0.33$) compared to the summary modality ($0.31$), though the difference is minimal. Similarly, the standard modality shows marginally better error metrics, with MSE ($1.65$ vs. $1.77$) and MAE ($0.95$ vs. $1.00$) indicating lower dispersion and deviation from the ground truth. Mode aggregation reflects a similar trend, with accuracy at $0.42$ for the standard modality and $0.38$ for the summary modality, alongside smaller errors (MSE: $1.81$ vs. $2.13$, MAE: $0.91$ vs. $0.98$). Median aggregation further supports this finding, with the standard modality showing slightly better performance across metrics (accuracy: $0.40$ vs. $0.38$, MSE: $1.80$ vs. $2.13$, MAE: $0.92$ vs. $0.93$). While errors in the summary modality are slightly more pronounced, both approaches achieve comparable accuracy levels. Bootstrap analysis confirms no statistically significant differences across accuracy and error measures. Overall, both modalities show similar effectiveness across several metrics, indicating that, despite the reduced volume of information provided to workers in the summary modality, it closely resembles the standard modality in terms of data quality. To test whether worker demographics influence the effectiveness of truthfulness assessments, we additionally run an Ordinary Least Squares (OLS) regression analysis [22]. The results show no statistically significant relationship between demographics and accuracy, indicating that these factors do not substantially impact assessment performance.
@@
Figure 3 shows the values of internal agreement across the two task modalities, both for the whole task (dotted lines) as well as for the single ground truth levels of the statements. For the Standard modality, the $\alpha$ values are generally lower than the Summary modality, reflecting overall less agreement among participants, with the only exception of the Pants On Fire truthfulness level. This suggests that when participants are presented with full-length content, there may be greater diversity in how the information is interpreted, possibly due to the complexity and volume of the provided information. Furthermore, the lowest $\alpha$ values appear in the categories in the middle of the scale, indicating that the claims with no extreme truthfulness score are also the most challenging for participants to consistently evaluate. Conversely, the Summary modality shows higher $\alpha$ values across all categories except Pants On Fire and for the overall task. This difference is statistically significant at the $p<0.05$ level, according to the Mann-Whitney U Test [34], indicating a higher degree of consistency among workers’ ratings. This improvement in internal agreement could be attributed to the condensed nature of the content, which might focus the workers evaluations and reduce both the cognitive load as well as the ambiguity or variability in interpretation that more extensive evidence might introduce. Notably, the alpha values are particularly higher in categories like Mostly False and True where the quality of the summarization may help clarify the key aspects of the statements.
Overall, these findings remark the impact of evidence summarization on the consistency of workers assessments. The higher internal agreement in the summary modality suggests that summarization not only maintains the essential content needed for the truthfulness assessment but may also aid in aligning worker scores, leading to more uniform interpretations and assessments of the truthfulness of statements.
@@
Table 3: Error types across truthfulness levels.
@@
Standard Pants On Fire 24 (24.0%) 76 (76.0%)
@@
Mostly False 15 (18.3%) 51 (62.2%) 34 (19.5%)
Half True 33 (48.5%) 41 (30.1%) 26 (20.6%)
Mostly True 51 (50.0%) 15 (14.7%) 34 (33.3%)
@@
Summary Pants On Fire 26 (26.0%) 74 (74.0%)
@@
Mostly False 12 (15.6%) 47 (61.0%) 41 (23.4%)
Half True 36 (42.9%) 34 (40.5%) 13 (16.6%)
Mostly True 40 (50.6%) 28 (32.0%) 32 (17.4%)
@@
A more fine-grained analysis, shown in the second part of Table 3, highlights the distribution of errors across modalities. In the Standard modality, overestimation emerges as the most frequent error type for categories such as Mostly False (62.2%). Conversely, correct assessments are most common in categories like Mostly True (50.0%) and Half True (48.5%), suggesting workers find it easier to correctly evaluate statements closer to the middle of the truthfulness spectrum. Underestimation occurs less frequently overall but is more pronounced for higher truthfulness categories. In the Summary modality, overestimation remains dominant in the same categories as in the Standard modality. However, correct assessments are slightly more evenly distributed, with the highest rates observed for Mostly True (50.6%) and Half True (42.9%). Overall, these findings suggest that both modalities exhibit similar patterns with respect to error types. We further analyze the mismatch magnitude for overestimation and underestimation errors across the Standard and Summary modalities and find similar error patterns. Both modalities exhibit comparable average mismatch values for overestimation (1.99) and slightly more negative values for underestimation in the Summary modality (-1.73 vs. -1.65). None of these differences are statistically significant, confirming the overall comparability of the two approaches.
@@
The results, shown in Figure 4, reveal a trend in which the accuracy of both modalities slightly improves, although without statistical significance, and the variance decreases as the number of workers per document increases. This pattern is consistent with expectations from crowdsourcing settings, where aggregating multiple assessments tends to mitigate individual biases and errors. Notably, the Summary modality achieves effectiveness levels comparable to the Standard modality across all sampling sizes. These indicate that summarized evidence can provide comparable results to full-length evidence in terms of both accuracy and consistency, and that this equivalence is achieved while potentially reducing the cognitive load on workers and the resources required for task completion.